Mary and Bryan together weigh 320 pounds. Bryan is 24 pounds heavier than Mary. How much does each person weigh
Found 2 solutions by macston, Miniml1:
Answer by macston(5194)

.
M=Mary's weight; B=Bryan's weight=M+24 lbs
M+B=320 lbs
M+(M+24 lbs)=320 lbs
M+M=296 lbs
2M=296 lbs
M=148 lbs
ANSWER 1: Mary weighs 148 pounds.
.
B=M+24lbs=148lbs+24lbs=172 lbs
ANSWER 2: Bryan weighs 172 pounds.
.
CHECK:
M+B=320lbs
148lbs+172lbs=320lbs
320lbs=320lbs

Answer by Miniml1(17)

You need to solve for x. 320 is on one side of the equation and Mary and Bryan's weight is on another side of the equation. Mary's weight is x and Bryan's weight is x+24 because he is 24 more pounds than Mary. The equation for this word problem is
320 = x + (x+24)
You can take everything out of the parenthesis, move 24 to the left side of the equation, and add the x's to make 2x
320 - 24 = 2x
When you move a number to the other side of the equation, its sign changes. In this case, positive 24 changed to negative 24. Now you can divide (320-24) by 2 because you want to make 2x to 1x. You have to divide because 2x indicates multiplication, and when you move anything to the other side of the equation, the sign changes.
296/2=x
x=148
Mary's age is 148, Bryan's age is 24 years older, 172